# SQ-GAN: Semantic Image Communications Using Masked Vector Quantization

### v1

@@ IV-A Data Augmentation
This section introduces a novel data augmentation technique, the semantic relevant classes enhancement, designed to solve the problem of underrepresented but semantically relevant classes in datasets. This new data augmentation technique is fundamentally different from previously proposed methods (e.g., [26, 27]). These approaches often focus on swapping existing objects – for example, replacing a "turn left" sign with a "stop" sign – or employ complex Neural Network (NN) architectures to introduce new objects into images. These approaches can lead to increased computational costs or fail to adequately address the under-representation of critical classes. In contrast, the proposed technique is straightforward, efficient, and fast.
@@
The process begins by collecting all "traffic signs" and "traffic lights" from the images and SSMs in the current mini-batch. For each image in the mini-batch, a random number $n$ between 00 and $25$ is selected, representing the number of objects to add. Then, $n$ objects are randomly chosen from the collected set and are carefully placed into the image and its SSM. The placement is done to avoid overlapping with existing instances of the same classes or with the other relevant class. An example of an augmented pair $(\mathbf{x},\mathbf{s})$ is represented in Fig. 6 with the original pair on the left and the data augmented version on the right. Other classic data augmentation techniques like cropping, rotation and color correction are also applied. However, it is interesting to notice the increase in "traffic signs" (yellow), and "traffic lights" (orange) in the augmented $\mathbf{x}$ and $\mathbf{s}$.
@@ IV-B Training $G_{\mathbf{s}}$
After data augmentation, the training of the sub-network $G_{\mathbf{s}}$ is implemented as depicted in Fig. 4, based on an adversarial approach VQ-GANs.
The input $\mathbf{s}$ and output $\hat{\mathbf{s}}$ tensors of $G_{\mathbf{s}}$ have shape $n_{c}\times H\times W$. We considered the following loss function:
$\mathcal{L}_{\text{SQ-GAN}}^{\mathbf{s}}=\mathcal{L}_{\text{WCE}}+\lambda_{\text{GAN}}\mathcal{L}_{\text{GAN}}+\lambda_{\text{vq}}\mathcal{L}_{\text{vq}}+\lambda_{\text{commit}}\mathcal{L}_{\text{commit}},$ (5)
where all the terms are defined as in [18], with the exception of the weighted cross-entropy loss $\mathcal{L}_{\text{WCE}}$, defined as:
@@
where $\mathbf{s}_{(h,w)}$ and $\hat{\mathbf{s}}{(h,w)}$ denote the one-hot encoded vectors at pixel location $(h,w)$ in $\mathbf{s}$ and $\hat{\mathbf{s}}$, respectively, and $\text{w}_{\mathbf{s}_{(h,w)}}\geq 0$ is the weight associated with the semantic class at that pixel. This choice is motivated by the nature of the SSM, which is a pixel-wise classification map where each pixel belongs to a specific semantic class. The weights are assigned to emphasize the importance of semantically relevant classes over non-relevant ones, encouraging the model to focus more on reconstructing the critical classes. Specifically, the weight factors are set as follows:
•   $\text{w}=1$ for the relevant classes "traffic signs" and
"traffic lights".
@@
•   $\text{w}=0.20$ for the non-relevant classes "sky" and
"vegetation".
@@
The other important component involved in training is the discriminator network $D_{disc}^{\mathbf{s}}$ [28]. This is composed of a convolutional layer, batch normalization layer and a leaky ReLU [29] repeated 3 times and followed by the last convolutional layer that maps the output to a single number. This value is the output of the discriminator used to classify the SSM as real or fake. During training, the masking fraction $m_{\mathbf{s}}$ is varied randomly, selected from a set of values ranging from 5% to 100% with and expected value of 35%. This approach allows the model to learn to compress the SSM at various levels of compression. The sub-network $G_{\mathbf{s}}$ is trained using the Adam optimizer [30] with a learning rate of $10^{-4}$ and a batch size of 8. The training is conducted for 200 epochs with early stopping to prevent overfitting.
@@ IV-C Training $G_{\mathbf{x}}$
The training of the sub-network $G_{\mathbf{x}}$ follows a similar approach. As said before, at this stage $G_{\mathbf{x}}$ is trained using the original image $\mathbf{x}$ and the original SSM $\mathbf{s}$, as illustrated in Fig. 5. Again, an adversarial approach typical of VQ-GANs is employed for training, with the loss function defined as:
$\mathcal{L}_{\text{SQ-GAN}}^{\mathbf{x}}=$ $\mathcal{L}_{WL_{2}}+\mathcal{L}_{\text{perc}}+\lambda_{\text{GAN}}\mathcal{L}_{\text{GAN}}+\lambda_{\text{vq}}\mathcal{L}_{\text{vq}}$ (7)
$+\lambda_{\text{commit}}\mathcal{L}_{\text{commit}},$ (8)
with the last three terms as in [18] and the first two described as follows. The weighted $\ell_{2}$ loss $\mathcal{L}_{WL_{2}}$, defined as:
@@
is designed to adjusts the importance of different semantic classes in the image. In (9), $\mathbf{x}_{(h,w)}$ and $\hat{\mathbf{x}}{(h,w)}$ are the pixel values at location $(h,w)$ of the tensors $\mathbf{x}$ and $\hat{\mathbf{x}}$, respectively, and $\text{w}_{\mathbf{s}_{(h,w)}}$ is the weight associated with the semantic class at that pixel, as given by the SSM $\mathbf{s}$. In our experiments, we used the following set of weights:
•   $\text{w}=1$ for the relevant classes "traffic signs" and
"traffic lights".
@@
•   $\text{w}=0$ for the non-relevant classes "sky" and
"vegetation".
@@
Among all possible choices of the weights, the most interesting is the one concerning the weights for "sky" and "vegetation", set to zero. This choice has been made to force the network to neglect the pixel-by-pixel reconstruction of these classes. In fact, the real color of the sky and trees in the image are semantically irrelevant.
The perceptual loss is instead the part that ensures that the reconstructed image maintains visual similarity to the original, preventing unrealistic alterations such as an unnatural sky color. To quantify this loss, we used Learned Perceptual Image Patch Similarity (LPIPS) [31], which evaluates the difference between two images in a latent space representation. The idea is that, if two images convey similar semantic content, like for example presence of relevant objects, as pedestrians, cars or traffic lights, each with the right shape, and well positioned relatively to each other, we should have a low (semantic) distortion, irrespective of the difference at the pixel level. LPIPS, instead of computing pixel-wise differences between $\mathbf{x}$ and $\hat{\mathbf{x}}$, compares the latent features extracted from the images by employing a deep neural network (DNN) pre-trained for object recognition. The DNN is somehow substituting a human to quantify the perceptual difference between two images. Formally, LPIPS is defined as:
$\mathcal{L}_{\text{perc}}=\sum_{l\in L}\frac{1}{H_{l}W_{l}}\sum_{(h_{j},w_{j})}\left\|\text{w}_{l}\odot\left(\phi_{i}(\mathbf{x})-\phi_{i}(\hat{\mathbf{x}})\right)\right\|_{2}^{2}$ (10)
where $\phi_{l}(\cdot)$ denotes the activation from the $l$-th layer of a pre-trained network $\phi$, $L$ is the set of layers used for feature extraction, $H_{l}$ and $W_{l}$ are the height and width of the feature map at layer $l$, and $\text{w}_{l}$ are learned weights that adjust the contribution of each layer. The operator $\odot$ denotes element-wise multiplication between the weight $\text{w}_{l}$ and all the elements of the difference $\left(\phi_{i}(\mathbf{x})-\phi_{i}(\hat{\mathbf{x}})\right)$ in any location coordinate $(h_{j},w_{j})$. By comparing the feature representations at multiple layers, LPIPSs captures perceptual differences at different scales and abstraction levels. Lower LPIPS values indicate better reconstruction quality.
@@
In recent years, several works have proposed ways to modify the discriminator by introducing various conditioning. For example, [32] conditioned the discriminator on the SSM to improve Semantic Segmentation (SSeg) retention, while [33] enforced the discriminator to focus on high-frequency components. However, these methods do not adequately address the issue at hand. To achieve the desired performance, it is essential to adjust the discriminator to minimize its focus on non-relevant regions. To this end, we make the following observations about the discriminator’s behavior:
@@
Based on these insights, a technique is proposed to reduce the discriminator’s focus on non-relevant semantic classes. The approach consists in artificially modifying the reconstructed image $\hat{\mathbf{x}}$ before it is evaluated by the discriminator. This modification aims to minimize the differences in non-relevant regions between $\mathbf{x}$ and $\hat{\mathbf{x}}$, bringing the data distribution of these regions of $\hat{\mathbf{x}}$ closer to the real distribution of $\mathbf{x}$. For example, if the generator reconstructs a tree with dark green leaves when in reality they are light green, the color in $\hat{\mathbf{x}}$ will be artificially shifted toward light green.
This artificial editing is performed by considering the residual between the real and reconstructed images, defined as $\mathbf{r}=\mathbf{x}-\hat{\mathbf{x}}$. By masking this pixel-wise difference between the two images and adding back a fraction of the residual to $\hat{\mathbf{x}}$ it is possible to obtain the new image $\hat{\mathbf{x}}_{rel}=\hat{\mathbf{x}}+\mathbf{w}_{rel}\odot\mathbf{r}$, where $\mathbf{w}_{rel}$ is the re-scaling relevance tensor. In our experiments, we have used the following values:
•   $\mathbf{w}_{rel}=0.90$ for the class "sky".
•   $\mathbf{w}_{rel}=0.80$ for the class "vegetation".
•   $\mathbf{w}_{rel}=0.40$ for the class "street".
•   $\mathbf{w}_{rel}=0$ for the other classes.
@@
The sub-network $G_{\mathbf{x}}$ is trained for different masking fractions $m_{\mathbf{x}}$. These values are selected from a finite set ranging from $5\%$ to $100\%$ with an expected value of $35\%$. The structure of $D_{disc}^{\mathbf{x}}$ is the same of $D_{disc}^{\mathbf{s}}$ and the training is performed using the Adam optimizer with a learning rate of $10^{-4}$ with batch size of $8$. The model is trained for $200$ epochs with early stopping.

### v2

@@ IV Training and Inference
In addition to the multi-step training approach, another crucial challenge is the identification and preservation of semantically relevant information. To address this problem, the training process has been further improved by proposing and incorporating a data augmentation step, adding semantically relevant classes, and a semantic-aware discriminator network. The idea is to emphasize the importance of semantically relevant classes. For example, in our numerical experiment we focused on dashboard camera views for assisted driving applications, where the semantic classes of interest are (for example) "traffic signs", "traffic lights", and "pedestrians". These objects typically occupy a small portion of the frame and are not always present in every frame. In contrast, non-relevant classes such as "sky", "vegetation", and "street", appear much more frequently and occupy a large portion of the images. Therefore, Data augmentation is used to ensure that frequent but task-irrelevant semantic classes do not dominate the training process.
This section is organized as follows. Section IV-A will present the data augmentation process as well as the proposed method for enhancing semantically relevant classes. Section IV-B will focus on the training of $G_{\mathbf{s}}$, while Section IV-C will focus on the training of $G_{\mathbf{x}}$ with the introduction of the proposed semantic-aware discriminator. Finally, Section IV-D will focus on the final fine-tuning process. The dataset used for training is the Cityscapes dataset [30] composed of 2975 pairs of images and associated SSMs.
IV-A Data Augmentation
This section introduces a novel data augmentation technique aimed at addressing the issue of underrepresented but semantically important classes in datasets. This new data augmentation technique is fundamentally different from previously proposed methods (e.g., [31, 32]). These approaches often focus on swapping existing objects – for example, replacing a "turn left" sign with a "stop" sign – or employ complex Neural Network (NN) architectures to introduce new objects into images. These approaches can lead to increased computational costs or fail to adequately address the under-representation of critical classes. In contrast, the proposed technique is straightforward, efficient, and fast.
@@ IV-A Data Augmentation
The process begins by collecting all "traffic signs" and "traffic lights" from the images and SSMs in the current mini-batch. For each image in the mini-batch, a random number $n$ between 0 and $25$ is selected, representing the number of objects to add. Then, $n$ objects are randomly selected from the collected set and carefully inserted into the image and its SSM. The placement is carried out in such a way as to avoid overlapping with existing instances of the same class or other semantically relevant classes. An example of an augmented pair $(\mathbf{x},\mathbf{s})$ is represented in Fig. 6 with the original pair on the left and the data augmented version on the right. Other classic data augmentation techniques like cropping, rotation and color correction are also applied. However, it is interesting to notice the increase in "traffic signs" (yellow), and "traffic lights" (orange) in the augmented $\mathbf{x}$ and $\mathbf{s}$.
@@ IV-B Training $G_{\mathbf{s}}$
The training of the sub-network $G_{\mathbf{s}}$ is implemented as depicted in Fig. 4, based on an adversarial approach VQ-GANs applied to the training set with data augmentation. The input and output tensors $\mathbf{s}$ and $\hat{\mathbf{s}}$ of $G_{\mathbf{s}}$ have a shape $n_{c}\times H\times W$. We considered the following loss function:
$\mathcal{L}_{\text{SQ-GAN}}^{\mathbf{s}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{\text{WCE}}+}$ (5)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hskip-11.38109pt+\lambda_{\text{GAN}}\left[\log D_{disc}^{\mathbf{s}}(\mathbf{s})+\log\left(1-D_{disc}^{\mathbf{s}}(G_{\mathbf{s}}(\mathbf{s}))\right)\right]}$ (6)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hskip-11.38109pt+\lambda_{\text{vq}}\|\text{sg}[\mathbf{z}^{{}^{\prime}\mathbf{s}}]-\mathbf{z}^{\mathbf{s}}_{q}\|_{2}^{2}+\lambda_{\text{commit}}\|\mathbf{z}^{{}^{\prime}\mathbf{s}}-\text{sg}[\mathbf{z}^{\mathbf{s}}_{q}]\|_{2}^{2}}$ (7)
$\hskip-17.07164pt:=\mathcal{L}_{\text{WCE}}+\lambda_{\text{GAN}}\mathcal{L}_{\text{GAN}}+\lambda_{\text{vq}}\mathcal{L}_{\text{vq}}+\lambda_{\text{commit}}\mathcal{L}_{\text{commit}}$ (8)
where the $\lambda$’s are hyperparameters that control the strength of each term. The last three terms in (8) are the same as in [20] for the conventional VQ-GAN. More specifically, $\mathcal{L}_{\text{GAN}}$ is the term responsible for adversarial training. This term encourages the generator to produce outputs that are indistinguishable from real semantic maps by penalizing the generator if the discriminator can differentiate between the real and generated images. The structure of the discriminator $D_{disc}^{\mathbf{s}}$ is that of a neural network as described in [33]. This is composed of a convolutional layer, batch normalization layer and a leaky ReLU [34] repeated 3 times and followed by the last convolutional layer that maps the output to a single number. This value is the output of the discriminator used to classify the SSM as real or fake. As a result, $\mathcal{L}_{\text{GAN}}$ improves the overall quality of the generated semantic reconstructions by guiding the model toward realistic results. The $\mathcal{L}_{\text{vq}}$ term is responsible for training the codebook used in the quantization process. It utilizes the $\text{sg}[\cdot]$, which is the stop-gradient operator. This operator prevents the gradients from being backpropagated to the encoder during training. By blocking the backpropagation, the encoder’s parameters are not updated, and the codebook is encouraged to adjust its codewords to better match the output of the encoder. The $\mathcal{L}_{\text{commit}}$ term ensures that the encoder does not overfit to only a small subset of the codebook entries. This is obtained by penalizing the difference between the encoder’s output $\mathbf{z}^{\prime}$ and the quantized codebook entry $\text{sg}[\mathbf{z}_{q}]$, thus promoting the encoder to use a more diverse set of codebook entries. As a result, the generalization and stability of the training process are improved while ensuring that the codebook entries are not updated during the training of the encoder.
The first term of the loss function in Eq. (8) is a novelty introduced in this work and consists of the weighted cross-entropy $\mathcal{L}_{\text{WCE}}$ defined as:
@@
where $\mathbf{s}_{(h,w)}$ and $\hat{\mathbf{s}}{(h,w)}$ denote the one-hot encoded vectors at pixel location $(h,w)$ in $\mathbf{s}$ and $\hat{\mathbf{s}}$, respectively, and $\text{w}_{\mathbf{s}_{(h,w)}}\geq 0$ is the weight associated with the semantic class in that pixel. This choice is motivated by the nature of the SSM, which is a pixel-wise classification map where each pixel belongs to a specific semantic class. The weights are assigned to emphasize the importance of semantically relevant classes over nonrelevant ones, encouraging the model to focus more on reconstructing the critical classes. Specifically, in the training of the model used in our numerical examples, these weights are set as follows:
•   $\text{w}=1$ for the relevant classes "traffic signs"
and "traffic lights".
@@
•   $\text{w}=0.20$ for the non-relevant classes "sky"
and "vegetation".
@@
After defining the loss function, the subnetwork $G_{\mathbf{s}}$ and the associate discriminator $D_{disc}^{\mathbf{s}}$, can be trained. During training, the masking fraction $m_{\mathbf{s}}$ is randomly varied, selected from a set of values ranging from 5% to 100% with an expected value of 35%. This approach allows the model to learn to compress the SSM at various levels of compression and increase the robustness of the model with respect to different conditions of the channel. If the task will require it, it is in fact possible to dynamically adapt the masking fraction without having to train another model from scratch, like other approaches like [12, 13] and [21] do. The subnetwork $G_{\mathbf{s}}$, and the associated discriminator $D_{disc}^{\mathbf{s}}$, are trained using the Adam optimizer [35] with a learning rate of $10^{-4}$ and a batch size of 8. The training is conducted for 200 epochs with early stopping to prevent overfitting.
@@ IV-C Training $G_{\mathbf{x}}$
The training of the subnetwork $G_{\mathbf{x}}$ follows a similar approach. As said before, at this stage $G_{\mathbf{x}}$ is trained using the original image $\mathbf{x}$ and the original SSM $\mathbf{s}$, as illustrated in Fig. 5. Again, an adversarial approach typical of VQ-GANs is employed for training, with the loss function defined as:
$\mathcal{L}_{\text{SQ-GAN}}^{\mathbf{x}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{Wl_{2}}+\mathcal{L}_{\text{perc}}+}$ (10)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hskip-17.07164pt+\lambda_{\text{GAN}}\left[\log D_{disc}^{\mathbf{x}}(\mathbf{x}|\mathbf{s})+\log\left(1-D_{disc}^{\mathbf{x}}(G_{\mathbf{x}}(\mathbf{x})|\mathbf{s})\right)\right]}$ (11)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hskip-17.07164pt+\lambda_{\text{vq}}\|\text{sg}[\mathbf{z}^{{}^{\prime}\mathbf{x}}]-\mathbf{z}^{\mathbf{x}}_{q}\|_{2}^{2}+\lambda_{\text{commit}}\|\mathbf{z}^{{}^{\prime}\mathbf{x}}-\text{sg}[\mathbf{z}^{\mathbf{x}}_{q}]\|_{2}^{2}}$ (12)
$\hskip-36.98857pt:=\mathcal{L}_{Wl_{2}}+\mathcal{L}_{\text{perc}}+\lambda_{\text{GAN}}\mathcal{L}_{\text{GAN}}+\lambda_{\text{vq}}\mathcal{L}_{\text{vq}}+\lambda_{\text{commit}}\mathcal{L}_{\text{commit}}$ (13)
where the last two terms and the $\lambda$’s are defined as in (8), while the first three terms are defined as follows. The weighted $\ell_{2}$ loss $\mathcal{L}_{Wl_{2}}$, defined as:
@@
is designed to adjust the importance of different semantic classes in the image. In (14), $\mathbf{x}_{(h,w)}$ and $\hat{\mathbf{x}}_{(h,w)}$ are the pixel values at location $(h,w)$ of the tensors $\mathbf{x}$ and $\hat{\mathbf{x}}$, respectively, and $\text{w}_{\mathbf{s}_{(h,w)}}$ is the weight associated with the semantic class at that pixel, as given by the SSM $\mathbf{s}$. Specifically, while training the model used in our numerical examples, we used the following weights:
•   $\text{w}=1$ for the relevant classes "traffic signs"
and "traffic lights".
@@
•   $\text{w}=0$ for the non-relevant classes "sky"
and "vegetation".
@@
Of particular interest is the zero weight assigned to the classes "sky" and "vegetation". This allows the network to neglect the pixel-by-pixel reconstruction of these classes. In fact, the real colors of the sky and trees for the task at hand are semantically irrelevant.
The perceptual loss $\mathcal{L}_{\text{perc}}$ is instead the part that ensures that the reconstructed image maintains visual similarity to the original, preventing unrealistic alterations such as an unnatural sky color. To quantify this loss, we used Learned Perceptual Image Patch Similarity (LPIPS) [36], which evaluates the difference between two images in a latent space representation. The underlying idea is that if two images convey similar semantic content, for example, if they both contain relevant objects such as pedestrians, cars, or traffic lights, each with the correct shape and positioned appropriately relative to one another, the semantic distortion should be low, regardless of any differences at the pixel level. LPIPS, instead of computing pixel-wise differences between $\mathbf{x}$ and $\hat{\mathbf{x}}$, compares the latent features extracted from the images by employing a deep neural network (DNN) pre-trained for object recognition. The DNN is somehow substituting a human to quantify the perceptual difference between two images. Formally, LPIPS is defined as:
$\mathcal{L}_{\text{perc}}=\sum_{l\in L}\frac{1}{H_{l}W_{l}}\sum_{(h_{j},w_{j})}\left\|\text{w}_{l}\odot\left(\phi_{l}(\mathbf{x})-\phi_{l}(\hat{\mathbf{x}})\right)\right\|_{2}^{2}$ (15)
[…]
@@
In recent years, several works have proposed ways to modify the discriminator by introducing various conditioning. For example, [37] conditioned the discriminator on the SSM to improve Semantic Segmentation (SSeg) retention, while [38] enforced the discriminator to focus on high frequency components. However, these methods do not adequately address the issue at hand. To achieve the desired performance, it is essential to adjust the discriminator to minimize its focus on non-relevant regions. To this end, we make the following observations about the discriminator’s behavior:
@@
Based on these insights, we propose to reduce the discriminator’s focus on non-relevant semantic classes. Our approach consists of artificially modifying the reconstructed image $\hat{\mathbf{x}}$ before it is evaluated by the discriminator and in (13) has been represented as conditioning the input of the discriminator as $D_{disc}^{\mathbf{x}}(\cdot|\mathbf{s})$. This modification aims to minimize the differences in non-relevant regions between $\mathbf{x}$ and $\hat{\mathbf{x}}$, bringing the data distribution of these regions of $\hat{\mathbf{x}}$ closer to the real distribution of $\mathbf{x}$. For example, if the generator reconstructs a tree with dark green leaves when in reality they are light green, the color in $\hat{\mathbf{x}}$ will be artificially shifted toward light green.
This artificial editing is performed by considering the residual between the real and reconstructed images, defined as $\mathbf{r}=\mathbf{x}-\hat{\mathbf{x}}$. By using the SSM to mask the pixel-wise difference between the two images, and adding back a fraction of the residual to $\hat{\mathbf{x}}$ it is possible to obtain the new image $\hat{\mathbf{x}}_{rel}=\hat{\mathbf{x}}+\mathbf{w}_{rel}\odot\mathbf{r}$. In this context $\mathbf{w}_{rel}$ represents the re-scaling relevance tensor that is obtained as a function of the SSM $\mathbf{s}$ and has the same shape as for the images, namely shape $3\times H\times W$. The element $(i,h,w)$ of the tensor $\mathbf{w}_{rel}$ is obtained from $\mathbf{s}$ by assigning the following values based on the class of the pixel $(h,w)$. In particular, in the training for our numerical examples, we used:
•   $\text{w}=0.90$ for the class "sky".
•   $\text{w}=0.80$ for the class "vegetation".
•   $\text{w}=0.40$ for the class "street".
•   $\text{w}=0$ for the other classes.
@@
The sub-network $G_{\mathbf{x}}$ is trained for different masking fractions $m_{\mathbf{x}}$, similarly to $G_{\mathbf{s}}$, to increase the robustness of the model with respect to different conditions of the channel. These values are selected from a finite set ranging from $5\%$ to $100\%$ with an expected value of $35\%$. The structure of $D_{disc}^{\mathbf{x}}$ is the same as $D_{disc}^{\mathbf{s}}$ and the training is performed using the Adam optimizer with a learning rate of $10^{-4}$ with batch size of $8$. The model is trained for $200$ epochs with early stopping.